OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0,+-6) and asymptotes at y=+-3/5x Please explain the steps!

OpenStudy (anonymous):

@satellite73

OpenStudy (amistre64):

can you start out with telling us the center, and the general setup for the hyperbolic equation?

OpenStudy (anonymous):

the center would have to be (0,0). and well it would have to be y-k^2/a^2 - x-h^2/b^2=1 @amistre64

OpenStudy (amistre64):

so far so good lets clean that up to be: (y/b)^2 - (x/a)^2 = 1 i like to keep my yb parts together and xa parts together, it helps with the asymptote information

OpenStudy (amistre64):

the slope of the asymptote is defined in as b/a we know b as the length from the center to the vertex in this case; so that should be 6 agreed?

OpenStudy (amistre64):



OpenStudy (amistre64):

from this we can determine that:\[\frac ba=\frac 35\] \[\frac 6a=\frac 35\] solve for a to finish the equation

OpenStudy (amistre64):

\[\frac{y^2}{b^2}-\frac{x^2}{a^2}=1\] \[\frac{y^2}{6^2}-\frac{x^2}{a^2}=1\]

OpenStudy (anonymous):

Can you explain how you got the 3/5? @amistre64

OpenStudy (amistre64):

i simply read the information in the problem ....

OpenStudy (anonymous):

Oh yes! sorry hahahahah. Thank you. Now I have answer choices, and the only answer choice it seems to fit would be y^2/36-x^2/100?

OpenStudy (amistre64):

well, since a = 10, id have to agree :)

OpenStudy (anonymous):

I had originally picked y^2/36-x^2/25 but i got it wrong

OpenStudy (amistre64):

you assumed the slope of the asymptote used exact values of b and a, instead of sclaed values

OpenStudy (amistre64):

*scaled values

OpenStudy (anonymous):

Thank you once again :D